### v1

@@ 3.4 Selection function
The full simulation of the SDC1 contains more than five million sources. Due to the simulated noise, only a fraction of these sources are detectable in the image. In Bonaldi et al. (2021), they estimated that only around 758000 sources are above the noise level by $5\sigma$ in the 560 MHz - 1000h image. This construction is problematic for supervised ML methods as they are very sensitive to wrong labeling during training. For an object detector, it goes two ways: i) if a source is detectable but not labeled as a target box, the network lowers the detection objectness of all predicted boxes that try to detect similar sources; ii) if there is a target box for a source that the network is incapable of detecting, it increases the objectness of non-representative features, likely background noise. Consequently, to achieve good detection performances, we must construct a training sample that is both complete and pure regarding the detectability of the sources it contains. For this, we defined a selection function based on the source properties of the provided True-catalog.
@@
Our selection function combines the apparent source flux $f_{a}$ in Jy, and a surface brightness estimate $S_{b}$ computed as
$S_{b}$ $=f_{a}/(W_{pix}\times H_{pix})\text{, where}$ (23)
$W_{pix}$ $=(2/P_{s})\times\max{(1.2,Bmaj)}\text{ and }$ (24)
$H_{pix}$ $=(2/P_{s})\times\max{(0.6,Bmin)}.$ (25)
@@
The size saturation values are in arcsec and correspond to a two and one-pixel size, respectively, and $P_{s}$ represents the pixel size in arcsec used to obtain the box size in pixels. The apparent source flux is obtained by multiplying the integrated source flux by the interpolated primary beam value at the central coordinates of the source. The selection function to keep a source in our training sample is then defined as
$\cases{S}_{b}>1.0\times 10^{-7}&\text{if }f>=7.0\times 10^{-6}\\ S_{b}>2.5\times 10^{-7}\text{ \& }f>1.65\times 10^{-6}\text{otherwise}.{}$ (26)
@@
We represent these cuts over a 2D histogram of the source surface against apparent flux for the training area in Fig. 7. In this figure, the top left patchy distribution corresponds to steep spectrum AGN. The flat spectrum AGN and star-forming galaxies follow the same principal distribution but are strongly unbalanced with only a few flat spectrum AGN. In practice, both types of AGNs are mostly undetectable with the 560 MHz - 1000h setup and, therefore, are strongly underrepresented after our selection function. This also contributes to the identified difficulty of the classification task in the SDC1 result paper (Sect. 5.3 in Bonaldi et al., 2021). The hard cut visible at the bottom left of the distribution results from the prefiltering done directly by SKAO, which is characterized by the ”selected” flag in the training catalog. Due to our size clipping, most of the surface range of this space is collapsed at a minimum surface, represented by the dotted grey line. We observed that lower clipping limits tend to increase the number of non-visible objects that pass the selection.
@@
We represent the effect of our selection function regarding the apparent flux distribution in Fig. 8. We also illustrate the effect of our selection on a small field in Fig. 9. We see that it misses some apparent compact signals, which can result from several effects: i) the local noise contribution can increase the perceived apparent flux of a very faint source, ii) it can correspond to blended faint point sources that add their flux at the same location, and iii) it can also be a bright, compact part of a globally large and faint source that does not respect the surface brightness limit. Regardless of their origin, these non-labeled compact signals will likely confuse the network training, but properly handling them would require a more complex selection function. We will see in Sects. 4.2 and 4.5 that the detector can still detect these types of sources, but since the perceived flux will not match the True catalog content, they are likely to be considered as false detections. We tried to adapt the selection process and our threshold values, but the current formulation produced the best results. We also tried to define our surface brightness using a more common astronomical size definition by convolving the $Bmaj$ and $Bmin$ with the synthesized beam. Still, it always resulted in lower detector scores after training.
In Appendix C.2, we discuss an alternative way to construct this function iteratively using the prediction of a naively trained network to evaluate the detectability of the sources.
@@
Figure 7: 2D histogram of the source surfaces as a function of their apparent flux. The surface is equivalent to $S_{b}$ but without minimum size clipping. The dotted grey line indicates the saturated minimum surface with clipping. The red dashed line represents our selection function.
@@
Figure 8: Source apparent flux distribution histogram in log scale and using log-bins for the True catalog of the training region and the training catalog after the selection function.
@@
Figure 9: Illustration of the selection function over a typical input field. The background image represents an identical $256{\times}256$ input patch in all frames, centered on $RA=0.1$ deg, $Dec=-30.2$ deg, using the re-normalized input dynamic but saturated at half the maximum value to increase visual contrast. The displayed grid corresponds to the detector output grid mapping for a network reduction factor of 16. The left frame is provided as an image reference, the middle frame represents all the target boxes from the True catalog, and the right frame represents the remaining target boxes after our selection function.
@@ 3.5 Training area
Using the challenge setup, the provided training area only spans a small part of the image center, corresponding to a surface of about $0.308$ square degree. The associated training catalog contains 190552 sources (using the pre-defined ”selected” flag), corresponding to roughly $3.6\%$ of the sources from the full True catalog. Applying our custom selection function to this region drops the number of sources to 33813. Due to the effect of the primary beam sensitivity over the image field, the central area is not a good representation of other parts of the image. The learned features and context awareness of a detector trained on this region are unlikely to generalize properly in other parts of the image. We represent the distribution of sources that pass our selection function based on apparent source brightness for the full image field using the full True cat in Fig. 10. The footprint of the primary beam sensitivity is clearly visible. The provided training area corresponds to the red box. A more suited training area definition could have been a narrow band over a full radius of the image, which we explore in Appendix C. To mitigate the generalization issue while still following the original challenge definition, two classical approaches would be to constrain the detector to be flux agnostic and only perform morphological detection or to reject all sources outside a given radius from the image center.
@@
We propose another approach that uses other regions of the image without adding any target sources. We observed that detecting an object above a given radius from the image center becomes very unlikely. From this, we added two ”noise only” regions to our training sample that are sufficiently far from the image center to be considered devoid of any detectable source. We selected two rectangular regions of identical width and height of 2000 and 5600 pixels, respectively, that are both vertically centered in the image but on opposite sides horizontally with a margin to the image edge of 250 pixels (Fig. 10). These regions lie between the primary beam’s main lobe and the first sidelobe. We acknowledge that the emptiness assumption is not completely true for multiple reasons and that it necessarily results in some non-labeled visible sources. Still, including cutouts from the ”mostly noise” context in the training sample is vastly beneficial regarding global detector performances. Training examples are mostly drawn randomly in the default training area but at a low $5\%$ rate in one of the two noise-only regions. This forces the network to understand the dynamic of parts of the image where it is not expected to detect anything. We observed that networks trained with this process can interpolate between the two regimes and provide much better results for the full image without manually excluding difficult regions. Compared to training on the default region only, the best achievable challenge score is increased by more than 10% with this approach.
@@
Figure 10: 2D histogram of the central coordinates of the sources from the full True catalog that pass our selection function. The red box indicates the default SDC1 training area, and the orange boxes indicate our additional noise-only regions.
@@ 3.6 Network backbone architecture
As introduced in Sect. 2.1, our method requires a fully convolutional neural network backbone to create a mapping from a 2D input image to a regular output grid. Efficient neural network architectures can be built by stacking convolutional layers while progressively reducing the spatial dimension to filter the relevant information and construct higher-level representations of the input content, forming a CNN (Simonyan & Zisserman, 2014; LeCun et al., 2015). With a fully convolutional structure, the last layer is responsible for encoding the output grid of the detector. Its spatial dimension represents the output grid, and the filters encode all the output vector elements for each grid cell. It also results in the output grid size being controlled by the reduction factor of the backbone network, which corresponds to the total spatial reduction of the network from its input to the output. The properties of a fully convolutional architecture and its impact on our method design are further detailed in Appendix A.1.
@@
With a fully convolutional architecture, we could technically use any input dimension (Appendix A.1) and even use different input sizes for training and prediction. In this section, we consider the input size to be $256{\times}256$ pixels (Sect. 3.8). Regarding the reduction factor, we found that a value of 16 resulted in the best detection accuracy, which means that the training output grid is composed of $16{\times}16$ cells. Considering the source density of the 560 MHz - 1000h image, each grid cell has to detect multiple sources, which was the motivation behind most of the adjustments we made to our association process in comparison to the classical YOLO (Appendix A.3).
@@
At this point, it would be tempting to adopt the classical YOLO-V2 architecture, darknet-19 (Redmon & Farhadi, 2016), as we successfully used it with CIANNA for other contexts with our custom association process (Appendix A.8). We discuss this possibility and explain why it would lead to poor performances in Appendix B. Instead, we meticulously explored increasingly complex custom backbone architectures. We specifically looked for an architecture that is both compute-efficient and capable of achieving high detection accuracy. An illustration of our final architecture is presented in Fig. 11.
@@
Our architecture is based on a few educated guesses, but it also requires exploration through score optimization. We first note that the spatial dimension is always reduced by convolution operations instead of pooling operations. In comparison to pooling operations, it helps preserve the apparent flux information, and it better represents continuous objects with no sharp edges. The first layer has larger filters to grasp more continuous luminosity profiles, while the second layer performs a local compression (both spatially and in the number of filters), mainly acting as a local noise filter. Then, for a few layers, we progressively increase the number of filters while decreasing the spatial dimension at a rate that maximizes computation performances. After the 7th layer, we start alternating large layers with $3{\times}3$ filters and smaller layers with $1{\times}1$ filters, which is typical of the YOLO darknet architecture. This structure alternates searches for local spatial coherency with representation compressions. It improves compute performance and reduces the global number of parameters compared to a more classical stacking of identically sized $3{\times}3$ layers. We also note that our third spatial-dimension reduction layer (8th layer in global) is also considered a compression layer. The second last layer has a 25% dropout rate, which is used for regularization (Srivastava et al., 2014).
@@
Following what is done in the darknet-19 architecture, we tested adding group normalization (GN, Wu & He, 2018) at various places in the network, but it almost always degraded the best achievable score by a few percent (Appendix B). This is likely because this type of normalization tends to lose the absolute values of the input pixels, making it more difficult to predict the flux accurately. Since the normalization is made on the full spatial dimension, it might also affect the dynamic in a way that makes faint sources more difficult to detect. The only place where GN produces a beneficial effect is near the end of the network, after the last spatial correlation, at a point where the flux value is likely fully re-encoded in the high-level features. This specific normalization layer has several beneficial effects, including a speed up and stabilization of the training process and also a small improvement of the best achievable score of about $2\%$.
@@
The full architecture contains 17 convolution layers for a total of around 12.62 million weights. We note that almost 76% of these weights are concentrated at the network’s end in the connections between layers 14 to 16. With this architecture, the receptive field for a given grid cell at the output layer is 100 pixels. This means that each detection unit will only capture information in an input region with a typical diameter of 100 pixels center on its grid cell. It limits both the maximum size of the sources that can be detected with this architecture and the context windows that each detection unit can use. We discuss the limits of the current network structure and potential architecture improvements in Sect. 5.2 and Appendix A.8.1.
@@
Figure 11: Illustration of our final CNN backbone architecture. The left column provides the layer structural properties, while the right column indicates the spatial output dimension for each layer starting from a $256{\times}256$ input size. The input image is on the top, and layers are stacked in order vertically. The width of a layer represents its number of filters. The green color indicates a layer that preserves the spatial dimension, while the red color indicates a reduction. The full output dimensions of the detector grid output are reminded at the end with the list of the predicted parameters for each box.
3.7 YOLO-CIANNA configuration
@@ 3.7.1 Input normalization
Input normalization is a critical requirement to obtain good detection performances in our specific context. This aspect is particularly difficult for astronomical images due to the very high dynamic range. Everyday-life images are usually encoded using three 8-bit integers, resulting in 256 possible values for each color. Converting an astronomical image to a similar format would induce a huge information loss. We found that 16-bit floating-point quantization was sufficient for the SDC1 image after other input transformations, but some astronomical datasets are likely to require 32-bit floating-point encoding. It is also useful to offset the raw dynamic to represent the low flux regime better. In our case, we redefine our minimum and maximum values as $min_{p}=4{\times}10^{-7}\,\mbox{Jy}.\mbox{beam}^{-1}$ and $max_{p}=4{\times}10^{-5}\,\mbox{Jy}.\mbox{beam}^{-1}$, and apply a scaled hyperbolic tangent for all pixel values. Our full renormalization can be summarized as
$p_{i}^{\prime}=\tanh{\left(3\times\frac{p_{i}-min_{p}}{max_{p}-min_{p}}\right)},$ (27)
where $p_{i}$ is the raw pixel value in $\mbox{Jy}.\mbox{beam}^{-1}$ clipped with the two limits, and $p_{i}^{\prime}$ is the pixel value as it is presented to the network. This normalization remaps all input data in the 0 to 1 range and grants most of this range to low signal values using an almost linear regime to maximize the available information to distinguish the sources from the noise. The counterpart is a flattening of the dynamic for high fluxes, but bright sources require less accuracy on the pixel values to obtain a good relative flux estimate.

### v2

@@ 3.4 Selection function
The full simulation of the SDC1 contains more than five million sources. Due to the added simulated noise, only a fraction of these sources are detectable in the image. In Bonaldi et al. (2021), they estimated that only around 758000 sources are above the noise level by $5\sigma$ in the 560 MHz - 1000h image. This construction is problematic for supervised ML methods as they are sensitive to wrong labeling during training. For an object detector, it goes two ways: i) if a source is detectable but not labeled as a target box, the network lowers the detection objectness of all predicted boxes that try to detect sources with similar features; ii) if there is a target box for a source that the network is incapable of detecting, it increases the objectness of nonrepresentative features, likely background noise. To achieve good detection performances, we must construct a training sample that is both complete and pure regarding the detectability of the sources it contains. For this, we defined a selection function based on the source properties of the provided True catalog.
@@
Our selection function combines the apparent source flux $f_{a}$ in Jy (Sect. 3.2), and the surface brightness $S_{b}$ computed as
$S_{b}$ $=f_{a}/(W_{\rm pix}\times H_{\rm pix})\text{, where}$ (23)
$W_{\rm pix}$ $=(2/P_{s})\times\max{(1.2,\textrm{Bmaj})}\text{ and }$ (24)
$H_{\rm pix}$ $=(2/P_{s})\times\max{(0.6,\textrm{Bmin})}.$ (25)
@@
The size saturation values are in arcsec and correspond to a two and one-pixel size, respectively, and $P_{s}$ represents the pixel size in arcsec used to obtain the box size in pixels. The selection function to keep a source in our training sample is then defined as
$\cases{S}_{b}>1.0\times 10^{-7}&\text{if }f_{a}>=7.0\times 10^{-6}\\ S_{b}>2.5\times 10^{-7}\text{ \& }f_{a}>1.65\times 10^{-6}\text{otherwise}.{}$ (26)
@@
The $f_{a}>1.65\times 10^{-6}$ hard cut in apparent flux is equivalent to a signal-to-noise-ratio (S/N) selection with $S/N>4.8$. However, we observed that S/N cuts were insufficient to achieve good results as too-low cuts failed to properly remove extended faint sources, while too-high cuts tended to remove detectable point sources. Consequently, we added the surface brightness cuts that helped to remove undetectable objects. We represent our selection cuts over a 2D histogram of the source surface against apparent flux for the training area in Fig. 7. In this figure, the top left patchy distribution corresponds to steep spectrum AGN. The flat spectrum AGN and star-forming galaxies distribute similarly and occupy the rest of the distribution, but they are strongly unbalanced with only a few flat spectrum AGN. In practice, both types of AGNs are mostly undetectable with the 560 MHz - 1000h setup, and are therefore strongly under-represented after our selection function. This also contributes to the identified difficulty of the classification task in the SDC1 result paper (Sect. 5.3 in Bonaldi et al., 2021). The hard cut at the bottom left of the distribution results from a per-class S/N prefiltering done directly by SKAO, characterized by the “selected” flag in the training catalog. Due to our size clipping, most of the surface range of this space is collapsed at a minimum surface, represented by the dotted gray line. We observed that lowering the clipping limits below the current values increases the number of nonvisible objects that pass our selection, which should be avoided.
@@
We represent the effect of our selection function on the source apparent flux distribution in Fig. 8. We also illustrate our selection on a small field in Fig. 9. We see that it misses some apparent compact signals, which can result from several effects: i) the local noise contribution can increase the perceived apparent flux of a faint source, ii) blended faint sources can add their flux at the same location, and iii) it can be a bright and compact part of an extended faint source below the surface brightness limit. Regardless of their origin, these nonlabeled compact signals will likely confuse the detector training. We tried to adapt the selection process and our threshold values, but the current formulation produced the best results. We also tried to define our surface brightness using a more common astronomical size definition by convolving the Bmaj and Bmin with the synthesized beam. Still, it resulted in lower detector scores after training.
We emphasize that this hand-made selection function relies on parameters from the True source catalog. Therefore, applying the method to observed data instead of simulated ones will require either using estimates from another approach to define the target or constructing a selection function that does not rely on these source properties (Sect. 5.3). In Appendix C.2, we discuss an alternative way to construct a selection function iteratively using the prediction of a naively trained detector to evaluate the detectability of the sources.
@@
Figure 7: Two-dimensional histogram of the source surfaces as a function of their apparent flux. The dotted gray line indicates the saturated minimum surface with clipping. The red dashed line represents our selection function, with all the sources on the right of the line being selected.
@@
Figure 8: Source apparent flux distribution histogram in log scale and using log-bins for the True catalog and the selected sources.
@@
Figure 9: Illustration of the selection function over a typical input field. The background image represents an identical $256{\times}256$ input patch in all frames, centered on $\textrm{RA}=0.1$ deg, $\textrm{Dec}=-30.2$ deg, using the renormalized input intensity but saturated at half the maximum value to increase visual contrast. The displayed grid corresponds to the detector output grid mapping for a network reduction factor of 16. The left frame is provided as an image reference, the middle frame represents all the target boxes from the True catalog, and the right frame represents the remaining target boxes after our selection function.
@@ 3.5 Training area
Using the challenge setup, the provided training area only spans a small part of the image center, corresponding to about $0.308$ square degree. The associated training catalog contains 190552 sources (using the predefined “selected” flag), corresponding to roughly $3.6\%$ of the sources from the full True catalog. Applying our custom selection function to this region drops the number of sources to 33813. Due to the effect of the primary beam sensitivity over the image field, the central area is not a good representation of other parts of the image. The learned features and context awareness of a detector trained on this region are unlikely to generalize properly to other parts of the image. We represent the distribution of sources that pass our selection function for the whole image field using the full True cat in Fig. 10. The footprint of the primary beam sensitivity is clearly visible. The provided training area corresponds to the red box. A more suited training area definition could have been a narrow band over a complete beam radius of the image field, which we explore in Appendix C. To mitigate the generalization issue while still following the original challenge definition, the detector could be constrained to either be flux agnostic and only perform morphological detection or to reject all sources outside a given radius from the image center.
@@
Instead, we propose another approach that uses other regions of the image without adding any target sources. We observed that detecting an object above a given radius from the image center becomes very unlikely. From this, we added two “noise only” regions to our training sample that are sufficiently far from the image center to be considered devoid of any detectable source. We selected two rectangular regions of identical width and height of 2000 and 5600 pixels, respectively, that are both vertically centered in the image but on opposite sides horizontally with a margin to the image edge of 250 pixels (Fig. 10). These regions lie between the main lobe and the first sidelobe of the primary beam. We note that the regions are, in fact, not fully empty and that it will result in a few unlabeled visible sources. Still, this approach remains vastly beneficial regarding global detector performances. During training, examples are drawn randomly in the default training area but also at a low $5\%$ rate in one of the two noise-only regions. This forces the detector to understand the input intensity dynamic of parts of the image where it is not expected to detect anything. We observed that detectors trained with this process can interpolate between the two regimes and provide much better results for the full image without manually excluding difficult regions. Other things being equal, the best achievable challenge score is increased by more than 10% with this approach compared to training only with the default region.
@@
Figure 10: Two-dimensional histogram of the central coordinates of the sources from the full True catalog that pass our selection function. The red box indicates the default SDC1 training area, and the orange boxes indicate our additional noise-only regions.
@@ 3.6 Network backbone architecture
As stated in Sect. 2.1, our method requires a fully convolutional neural network backbone to create a mapping from a 2D input image to a regular output grid. Efficient neural network architectures can be built by stacking convolutional layers while progressively reducing the spatial dimension to filter the relevant information and construct higher-level representations of the input content (Simonyan & Zisserman, 2015; LeCun et al., 2015). With a fully convolutional structure, the last layer is responsible for encoding the output grid of the detector. Its spatial dimension represents the output grid, and the filters encode all the output vector elements for each grid cell. It also results in the output grid size being controlled by the reduction factor of the backbone network, which corresponds to the total spatial reduction of the network from its input to the output. The properties of a fully convolutional architecture and its impact on our method design are further detailed in Appendix A.1.
@@
With a fully convolutional architecture, the input size does not impact the model structure and number of parameters. In this section, we consider the input size to be $256{\times}256$ pixels (Sect. 3.8). Regarding the reduction factor, we found that a value of 16 resulted in the best detection accuracy, which means that the training output grid is composed of $16{\times}16$ cells. Considering the source density of the 560 MHz - 1000h image, each grid cell has to detect multiple sources, which was the motivation for using a prediction-aware association process (Appendix A.3).
@@
At this point, it would be tempting to adopt the classical darknet-19 backbone network introduced in YOLO-V2 (Redmon & Farhadi, 2017). We notably used it successfully with our custom association process for other contexts (Appendix A.8). We discuss this possibility and explain why it would lead to poor performances in Appendix B. Instead, we meticulously explored increasingly complex custom backbone architectures. We specifically looked for an architecture that is both computationally efficient and capable of high detection accuracy. An illustration of our final architecture is presented in Fig. 11.
@@
Our final architecture was based on a few educated guesses, but it also required exploration through score optimization. We note that the spatial dimension is always reduced by convolution operations instead of pooling operations, which helps preserve the apparent flux information and better represents continuous objects with no sharp edges. The first layer has larger filters to extract continuous luminosity profiles better. The second layer performs a local compression (both spatially and in the number of filters), mainly acting as a local noise filter. Then, for a few layers, we progressively increase the number of filters while decreasing the spatial dimension at a rate that maximizes computing efficiency. Starting with the seventh layer, we begin alternating large layers with $3{\times}3$ filters and smaller layers with $1{\times}1$ filters, which is typical of the YOLO darknet architecture. This structure alternates searches for local spatial coherency with representation compressions. It improves compute performance and reduces the global number of parameters compared to a more classical stacking of identically sized $3{\times}3$ layers. We note that our third spatial-dimension reduction layer (eighth layer in global) is also considered a compression layer in this scheme. The second last layer has a 25% dropout rate, which is used for regularization (Srivastava et al., 2014).
@@
We tested adding group normalization (Wu & He, 2018) at various places in the network, but it almost always degraded the best achievable score by a few percent (Appendix B). This is likely because in-network normalization tends to lose the absolute values of the input pixels, making it more difficult to predict the flux accurately. Since group normalization works at the scale of the whole spatial dimension, it might also affect the input dynamic in a way that makes faint sources more challenging to detect in the presence of bright sources in the image. The only place where it produces a beneficial effect is near the end of the network, after the last spatial correlation, where the flux value is likely fully re-encoded in the high-level features. This specific normalization layer has several beneficial effects, including a speed up and stabilization of the training process and a slight improvement of the best achievable score of about $2\%$.
@@
The complete architecture contains 17 convolution layers for around 12.62 million weights, but 76% of these weights are concentrated at the end of the network in the connections between layers 14 to 16. With this architecture, the receptive field for a given grid cell at the output layer is about 100 pixels. It limits both the maximum size of the sources that can be detected with this architecture and the context windows accessible to each detection unit. We discuss the limits of the current network structure and potential architecture improvements in Sect. 5.2.
@@
Figure 11: Illustration of our final CNN backbone architecture. The left column provides the layer structural properties, while the right column indicates the spatial output dimension for each layer starting from a $256{\times}256$ input size. The input image is on the top, and layers are stacked in order vertically. The width of a layer represents its number of filters. The green color indicates a layer that preserves the spatial dimension, while the red indicates a reduction. The output grid dimension and the list of the predicted parameters for each box are also indicated.
3.7 YOLO-CIANNA configuration for the SDC1
@@ 3.7.1 Input normalization
Input normalization is critical to obtain good detection performances in our specific context. This aspect is key for astronomical images due to their very high dynamic range. Everyday-life images are usually encoded using three 8-bit integers, resulting in 256 possible values for each color. Converting an astronomical image to a similar format often induces a significant information loss. We note that some off-the-shelves deep learning detectors or frameworks implicitly convert images to this format without warning, sometimes explaining poor performances. We found that 16-bit floating-point quantization was sufficient for the SDC1 image after other input transformations, but some astronomical datasets are likely to require 32-bit floating-point encoding. It is also useful to offset the raw dynamic to better represent the low flux regime. In our case, we redefine our minimum and maximum values as $\min_{p}=4{\times}10^{-7}\,\mbox{Jy}\,\mbox{beam}^{-1}$ and $\max_{p}=4{\times}10^{-5}\,\mbox{Jy}\,\mbox{beam}^{-1}$, and apply a scaled hyperbolic tangent for all pixel values, which can be summarized as
$p_{i}^{\prime}=\tanh{\left(3\times\frac{p_{i}-\min_{p}}{\max_{p}-\min_{p}}\right)},$ (27)
where $p_{i}$ is the raw pixel value in $\mbox{Jy}\,\mbox{beam}^{-1}$ clipped with the two limits, and $p_{i}^{\prime}$ is the pixel value as it is presented to the detector. This normalization remaps all input data in the 0 to 1 range and grants most of this range to low signal values using an almost linear regime to help distinguish faint sources from the noise. The counterpart is a flattening of the dynamic for high fluxes, but bright sources require less accuracy on the pixel values to obtain a good relative flux estimate.